40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 9 01:29:26 2017, gautam, Update, LSC, DRMI Nosie Budget v3.0 7x
    Reply  Mon Oct 9 11:55:01 2017, Koji, Update, LSC, DRMI Nosie Budget v3.0 
       Reply  Mon Oct 9 22:18:34 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_dark.pdf
          Reply  Tue Oct 10 22:04:06 2017, rana, Update, LSC, AS55Q Dark Noise 
          Reply  Wed Oct 11 14:42:03 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
             Reply  Wed Oct 11 19:31:32 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
                Reply  Thu Oct 12 01:50:11 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_dark.png
                   Reply  Thu Oct 12 12:17:28 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_AS55_40m.pdf
                      Reply  Fri Oct 13 12:26:12 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_DemodNoises.pdf
                         Reply  Mon Oct 16 16:01:04 2017, gautam, Update, LSC, AS55Q Dark Noise DemodBoardwOscope.pdfDemodBoardwOscope_ASD.pdf
                            Reply  Tue Oct 17 19:31:53 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_TFs.pdfAS55_DemodNoises.pdf
    Reply  Tue Nov 7 17:45:05 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 SRCL_MICH_whitenoise_tf.pdfPRCL_MICH_whitenoise_tf.pdfMICH_aux.pdfC1NB_disp_40m_MICH_NB_2017-10-08.pdf
       Reply  Tue Nov 7 22:56:21 2017, gautam, Update, LSC, DRMI locking recovered DRMI_7Nov20178.png
          Reply  Wed Nov 8 00:28:16 2017, gautam, Update, LSC, Laser intensity coupling measurement attempt PSL_RIN.pdf
       Reply  Wed Nov 8 09:37:45 2017, rana, Update, LSC, DRMI Nosie Budget v3.1 
          Reply  Wed Nov 8 09:59:12 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 
             Reply  Wed Nov 15 01:37:07 2017, gautam, Update, LSC, DRMI low freq. nosie improved C1NB_disp_40m_MICH_NB_2017-11-15.pdfOL_A2L_couplings.pdf
                Reply  Thu Nov 16 00:53:26 2017, gautam, Update, LSC, DRMI noise sub-budgets C1NB_a2l_40m_MICH_NB_2017-11-15.pdfC1NB_aux_40m_MICH_NB_2017-11-15.pdf
Message ID: 13372     Entry time: Wed Oct 11 14:42:03 2017     In reply to: 13369     Reply to this: 13374
Author: gautam 
Type: Update 
Category: LSC 
Subject: AS55Q Dark Noise 

I keep adding traces to this plot, here is the most complete one I have now. Looks like the input noise to the D040179 (measured at "Q out" SMA jack of D990511 with RFPD input terminated) is ~10nV/rtHz. This supports the hypothesis that something is wonky on the daughter board, because the purple trace should only be the quad sum of the orange and green traces. I will pull it out and have a look.

Some other follow-ups on the questions raised at the meeting:

  1. Doesn't look like I've implemented thin film resistors on the input of the coil driver boards. De-whitening boards have the critical signal path resistors (judged as the ones with largest contribution as per LISO model) changed to thin film. Pictures are here.
  2. I think I didn't make a full elog of my demod board efficiency investigations, but from my notes and Attachment #4 of elog 12972, I calculated the gain in the signal path as the ratio of Vpp_out / Vpp_in.
Quote:

I measured the output voltage noise of the Q output of the AS55 Demod Board with the PSL shutter closed, using the SR785 (see Attachment #1). The measured noise is consistent with the expected number of ~120nV/rtHz around 100Hz. I had measured the gain of this board from RFPD input to Q output to be ~5.1: so if the PD dark noise is 16nV/rtHz, this would be amplified to ~80nV/rtHz. Still a discrepancy of ~50%. I didn't measure the noise with the PD input terminated. Added the noise of the demod board output with the RFPD input terminated. The level of ~100nV/rtHz seems consistent with the actual PD dark noise being ~80nV/rtHz, as their quadrature sum is around 130nV/rtHz. Need to dig up the schematics for the demod board + daughter board, and check against LISO, to see if this is consistent with what is expected.

Also - I think I was using the wrong value of the DC power on the AS55 photodiode for shot noise calculations - 13mW was for REFL55, not AS55. I did a crude measurement of the power by sticking the Ophir power meter (filter removed) in front of the AS55 PD with the Michelson flashing around, and noticed the maximum value registered was ~1.2mW. So in the DRMI lock, there would be ~2.4mW, which is 10x lower than the value I was assuming. I've made the correction in the NB code, for the next time the plot is generated. A more rigorous measurement would involve sticking the Ophir in front of the AS110 PD during a DRMI lock. The light from the AS port is split by a 50-50 BS to the AS55 and AS110 PDs (so measuring at AS110 is a reasonable proxy for power at AS55), and the AS110 signals are not used for triggering in the DRMI lock, so this is feasible.

 

 

Attachment 1: AS55Q_darkNoises.pdf  167 kB  | Show | Show all
ELOG V3.1.3-